OpenStudy (annie12m):

Can you figure out how many questions I need to get correct to past? ~80% or up to past ~300 points in total ~5 points per question ~60 questions in total

OpenStudy (igreen):

\(\sf \dfrac{80}{100} = \dfrac{x}{300}\) Solve for 'x', can you do that?

OpenStudy (igreen):

Well, the common denominator is 300, and 300 / 100 is 3..so you can just multiply that to 80.

OpenStudy (anonymous):

oh that is so easy

OpenStudy (annie12m):

3 * 80? = 240

OpenStudy (igreen):

Correct..so you need at least 240 points to pass. Since each question is worth 5 points, divide it by 5.

OpenStudy (annie12m):

240 / 5 = 4

OpenStudy (annie12m):

48

OpenStudy (annie12m):

so 48 questions?

OpenStudy (igreen):

Yes..that's our answer.

OpenStudy (annie12m):

thanks so much!

OpenStudy (anonymous):

0.8 ( 60 ) = 48

OpenStudy (igreen):

You need to get 48 out of the 60 questions correct to have an 80%.

OpenStudy (igreen):

No problem.